Helen and James had between them 160 CDs. Helen said to James, ''if i had 5 more CDs, I would have twice as many as you''. How many CDs do they have each? Let the number of CDs that James has be x

Answer by ikleyn(52798)

.
Helen and James had between them 160 CDs.
Helen said to James, ''if i had 5 more CDs, I would have twice as many as you''.
How many CDs do they have each? Let the number of CDs that James has be x
~~~~~~~~~~~~

James = x CDs

Helen = 160-x  CDs.


Write an equation as you read the problem


    (160-x) + 5 = 2x.


Simplify and find x


    160+5 = 2x + x

    165   = 3x

     x    = 165/3 = 55.


ANSWER.  James had 55 CDs;  Helen had (160-55) = 105 CDs.

Solved.